# Secure Multi-Key Homomorphic Encryption with Application to Privacy-Preserving Federated Learning

### v1

@@ Abstract
Multi-Key Homomorphic Encryption (MKHE), proposed by López-Alt et al. (STOC 2012), allows for performing arithmetic computations directly on ciphertexts encrypted under distinct keys. Subsequent works by Chen and Dai et al. (CCS 2019) and Kim and Song et al. (CCS 2023) extended this concept by proposing multi-key BFV/CKKS variants, referred to as the CDKS scheme. These variants incorporate asymptotically optimal techniques to facilitate secure computation across multiple data providers. In this paper, we identify a critical security vulnerability in the CDKS scheme when applied to multiparty secure computation tasks, such as privacy-preserving federated learning (PPFL). In particular, we show that CDKS may inadvertently leak plaintext information from one party to others. To mitigate this issue, we propose a new scheme, SMHE (Secure Multi-Key Homomorphic Encryption), which incorporates a novel masking mechanism into the multi-key BFV and CKKS frameworks to ensure that plaintexts remain confidential throughout the computation. We implement a PPFL application using SMHE and demonstrate that it provides significantly improved security with only a modest overhead in homomorphic evaluation. For instance, our PPFL model based on multi-key CKKS incurs less than a 2×2$\times$2 × runtime and communication traffic increase compared to the CDKS-based PPFL model. The code is publicly available at https://github.com/JiahuiWu2022/SMHE.git.
Index Terms:
Multi-key homomorphic encryption, masking scheme, privacy protection, federated learning.
I Introduction
@@
Homomorphic encryption (HE) is a cryptographic technique that enables computations to be performed directly on encrypted data, eliminating the need for decryption during the process. This property allows for the secure processing of sensitive data while preserving its confidentiality. For decades, constructing a fully homomorphic encryption (FHE) scheme, capable of supporting arbitrary computations on ciphertexts, remained an open problem until Gentry’s groundbreaking work [1]. Since then, significant advancements have been achieved in the field, leading to various HE schemes such as BFV [2, 3], GSW [4], BGV [5], TFHE [6], and CKKS [7]. These schemes have expanded the practicality and efficiency of HE, making it a vital tool for secure computations in modern applications. One of the key features of HE is its ability to enable secure computation “on-the-fly,” meaning that data owners are not required to be actively involved during the computation process. Instead, the evaluation can be carried out entirely by a public server. This makes HE particularly appealing for scenarios such as cloud computing, where data is processed remotely, and privacy preservation is of paramount importance.
@@
In recent years, the demand for secure multiparty computation (MPC) protocols has surged, driven by applications such as federated learning [8]. MPC allows multiple parties to collaboratively evaluate a function or circuit on their private inputs without revealing any information beyond the final result. However, traditional single-key HE schemes are not well-suited for such multi-party settings. A major limitation arises when multiple data sources are involved, as standard HE schemes typically require all data to be encrypted under the same encryption key. This requirement grants the entity possessing the corresponding decryption key full access to the encrypted data, which raises issues regarding privacy and the reliance on centralized trust.
@@
In THE, multiple parties collaboratively generate a common public key, with the corresponding secret key shared among them through a secret-sharing mechanism. Although THE schemes generally achieve performance levels comparable to single-key HE and are typically more efficient than multi-key HE schemes, they suffer from a critical limitation: reliance on a static key access structure. In other words, all participants must be predetermined and fixed during the initial setup phase. In contrast, this work focuses on MKHE, which offers significant advantages in terms of flexibility and reduced interaction requirements. Specifically, MKHE schemes allow each participant to independently generate their own public-secret key pairs without needing knowledge of other parties’ keys. This property enables operations on ciphertexts encrypted under different keys and allows computations to be performed in a public cloud environment without establishing a common public key. Moreover, recent advances in MKHE have introduced full dynamism, allowing computations to be performed on multi-key ciphertexts without predefined circuits. Arbitrary circuits can be evaluated in real time, and new participants or ciphertexts can be incorporated into ongoing evaluations at any stage. This flexibility facilitates the construction of SMC protocols on top of MKHE, leveraging its dynamic and adaptive nature [17]. The ability to seamlessly integrate new participants and ciphertexts on-the-fly is a key advantage, making MKHE particularly well-suited for applications such as federated learning, where the set of participants may change throughout the computation process.
@@
I-A Challenges and Contributions
Recent MKHE schemes, particularly the CDKS constructions [20, 21], have significantly improved practicality. However, they exhibit security vulnerabilities in real-world multiparty computation scenarios. To address this, we revisit the security assumptions of CDKS-style constructions and identify several critical challenges. Our main contributions are summarized below.
Challenge 1: Inadequate Security Guarantees in CDKS for MPC. The CDKS schemes [20, 21] lack adequate security guarantees in MPC settings. In these schemes, ciphertexts from different parties are expanded into a unified format embedding all parties’ identity information for joint evaluation. Our analysis reveals that this expansion and evaluation process may inadvertently leak plaintext data. A detailed examination of this vulnerability, rooted in the ciphertext expansion mechanism, is provided in Section IV-B.
Our Solution: We propose a secure multi-key HE framework named SMHE. Our approach introduces a new ciphertext expansion and evaluation method by leveraging masking primitives, originally developed in GSW-based MKHE [16, 17], and we adapt them to the polynomial-based structures of BFV and CKKS.
At a high level, a masking refers to an encryption of a random value that is added to a fresh ciphertext. Since the masking is itself a semantically secure ciphertext, it reveals no information about the random value it encodes. This encrypted randomness serves as a one-time pad to conceal the participant’s plaintext during ciphertext expansion and intermediate homomorphic evaluation. This effectively mitigates the attack we later demonstrate against CDKS, in which adversaries exploit partial decryption results to infer individual plaintexts.
The core challenge lies in efficiently constructing such encrypted masks that can be correctly removed during final decryption to preserve result correctness, without undermining the inherent dynamism of the MKHE framework. In GSW-based MKHE schemes [16, 17], the masking scheme allows each party $P$ to produce auxiliary information $U$ during encryption. This information, published together with the ciphertext $C$, leaks no information about the underlying plaintext and is used to support joint computation. When a second participant $P^{\prime}$ joins with public key $pk^{\prime}$, the auxiliary information $U$ and $pk^{\prime}$ can be used to construct a masking matrix $X$ satisfying
$\begin{split}(sk,sk^{\prime})\left(\begin{array}[]{cc}C&X\\ 0&C\\ \end{array}\right)&=(skC,skX+sk^{\prime}C)\\ \approx(\mu\cdot sk\Theta,\mu\cdot sk^{\prime}\Theta)&=\mu(sk,sk^{\prime})\Theta,\end{split}$ (1)
where $sk$ and $sk^{\prime}$ are the secret keys of the two parties, $\mu$ is the plaintext of the party $P$, and $\Theta$ is a fixed public matrix. Define the concatenated secret key $\overline{sk}=(sk,sk^{\prime})$ and the expanded ciphertext $\overline{C}=\left(\begin{array}[]{cc}C&X\\ 0&C\\ \end{array}\right)$. Then, using GSW decryption, the plaintext is recovered as $\mu\overline{sk}\Theta\cdot\Theta^{-1}W\approx\mu$, where $W$ is a fixed GSW decryption vector. Similarly, for party $P^{\prime}$, with auxiliary information $U^{\prime}$ and ciphertext $C^{\prime}$, the expansion yields: $\overline{sk}\left(\begin{array}[]{cc}C^{\prime}&0\\ X^{\prime}&C^{\prime}\\ \end{array}\right)=(skC^{\prime}+sk^{\prime}X^{\prime},sk^{\prime}C^{\prime})\approx\mu^{\prime}\cdot\overline{sk}\Theta,$ where $X^{\prime}$ is constructed from $U^{\prime}$ and $pk$, and $\mu^{\prime}$ is the plaintext of $P^{\prime}$. The homomorphic addition of the two parties’ expanded ciphertexts is given by: $\overline{C}+\overline{C}^{\prime}=\left(\begin{array}[]{cc}C+C^{\prime}&X\\ X^{\prime}&C+C^{\prime}\\ \end{array}\right),$ and its decryption yields:
$\overline{sk}\left(\begin{array}[]{cc}C+C^{\prime}&X\\ X^{\prime}&C+C^{\prime}\\ \end{array}\right)\cdot\Theta^{-1}W$
=\ = $\big{(}sk(C+C^{\prime})+sk^{\prime}X^{\prime},skX+sk^{\prime}(C+C^{\prime})\big{)}\cdot\Theta^{-1}W$
@@ I-A Challenges and Contributions
This ensures the correctness of homomorphic addition. Homomorphic multiplication can be handled similarly, as shown in [16, 17]. By using this construction, ciphertexts encrypted under different keys are transformed into a unified expansion format, enabling efficient homomorphic operations in a dynamic multi-key setting. This allows each party to encrypt data independently while still enabling collaborative computation, without predetermining key structures or decryption.
Challenge 2: While the masking scheme proposed in [16, 17] effectively ensures the security and the correctness of GSW-based MKHE schemes, it cannot be directly applied to CKKS or BFV schemes for the following reasons:
R1: GSW ciphertexts are represented as matrices $C\in\mathbb{Z}_{Q}^{n\times m}$, whereas CKKS/BFV ciphertexts are polynomial pairs $ct=(c_{0},c_{1})\in R_{Q}^{2}$. This structural difference makes it challenging to construct a multi-key CKKS/BFV joint decryption equation of the form
@@
(where $cx$ is a masking with the same form as $ct$ constructed by a auxiliary information we denote as $\Gamma$, and $\theta$ is a constant term), analogous to the GSW multi-key decryption equation $\big{(}sk(C+X)+sk^{\prime}C\big{)}\cdot\Theta^{-1}W\approx\mu\cdot\overline{sk}\cdot W=\mu$.
R2: The GSW masking scheme requires large extended ciphertexts for correct demasking, with each single-bit encryption resulting in a ciphertext of size $O(N^{2})$, quadratic in the number of parties. This leads to significant computational and communication overhead. Therefore, it is necessary to design a new masking scheme for CKKS/BFV that ensures the expanded ciphertext size scales linearly with $N$, thereby improving both computational and space efficiency.
Our Solution: To address the challenges outlined in R1 and R2, we propose a novel masking scheme specifically designed for CKKS/BFV. To ensure the correctness of the joint decryption of the masked ciphertext (implicitly implying the unmasking operation followed by decryption), we add an additional “demasking” ciphertext to the masked ciphertext. Specifically, we construct a masking ciphertext $cx=(x_{0},x_{1})$ and a “demasking” ciphertext $cz=(z_{0},z_{1})$, where $cx$ contains a random mask $r$. The condition $\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\approx 0$ must hold, and the correct joint decryption is then given by:
$\qquad\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle+\langle sk,ct\rangle\approx\mu$
$\Rightarrow\langle(1,s,s^{\prime}),(z_{0}+x_{0}+c_{0},x_{1}+c_{1},z_{1})\rangle\triangleq\langle\overline{sk},\overline{ct}\rangle\approx\mu$
where $\overline{sk}=(1,s,s^{\prime})$ is the concatenated secret key and $\overline{ct}=(z_{0}+x_{0}+c_{0},x_{1}+c_{1},z_{1})\triangleq(\bar{c}_{0},\bar{c}_{1},\bar{c}_{2})$ is the expanded ciphertext. Notably, the size of our expanded ciphertext $\overline{ct}$ is linear with the number of participants. The remaining task is to construct the equation
@@
Taking the CKKS scheme as an example (which is similar to the BFV scheme), encryption of a plaintext $\mu$ is expressed as $ct\leftarrow\texttt{Encrypt}(pk,\mu)=w\cdot pk+(\mu+e_{0},e_{1})\pmod{Q}$, where $w$ is a random polynomial, $sk=(1,s)$ is the secret key, $pk=(b=-as+e,a)$ is the public key, $a$ is a polynomial in a given ring $R_{Q}$, and $e,e_{0},e_{1}$ represent small errors. The construction of Eq. (2) is then described as follows.
Let $cz$ be a ciphertext representing zero. It then follows that:
$\langle sk^{\prime},cz\rangle=\langle(1,s^{\prime}),r\cdot pk+(e_{0},e_{1})\rangle=r(b-b^{\prime})+e_{z}\pmod{Q},$
where $r$ is a random polynomial, $b^{\prime}=-as^{\prime}+e^{\prime}$, and $cz\leftarrow\texttt{Encrypt}(pk,0)=r\cdot pk+(e_{r_{0}},e_{r_{1}})\pmod{Q}$ represents the encryption of the plaintext zero, with $e^{\prime},e_{r_{0}},e_{r_{1}}$ being small errors. The masking is performed using the random polynomial $r$, which is encrypted as auxiliary information $\Gamma\leftarrow\texttt{Encrypt}(pk,r)=w\cdot pk+(r+e_{x_{0}},e_{x_{1}})$, where $e_{x_{0}},e_{x_{1}}$ are small errors. The evaluator then sets $cx=\Gamma(b^{\prime}-b)$ to make $\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\approx r(b^{\prime}-b)+r(b-b^{\prime})=0$.
Challenge 3: In the decryption process, the evaluator is expected to compute a correction term $cx=\Gamma(b^{\prime}-b)$ such that $\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\approx 0.$ However, since $b^{\prime}-b\in R_{Q}$ is essentially a random polynomial, we have:
@@
$=(b^{\prime}-b)(r+we+e_{x_{0}}+e_{x_{1}}s)=r(b^{\prime}-b)+E\pmod{Q}$
where $E=(b^{\prime}-b)(we+e_{x_{0}}+e_{x_{1}}s)\in R_{Q}$ is a large and non-negligible noise term. As a result, the desired equation $\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\approx 0$ fails to hold due to this accumulated noise, rendering the decryption equation invalid.
Solution: To control the noise growth, we apply the gadget decomposition technique to reduce the growth of $E$, ensuring that $\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\approx 0$. For a detailed explanation of the gadget decomposition technique, the reader is referred to Section III-C, and for a comprehensive description of our masking scheme, Section V-A provides further details.
II Related Work

### v2

@@ Abstract
Multi-Key Homomorphic Encryption (MKHE), proposed by López-Alt et al. (STOC 2012), allows for performing arithmetic computations directly on ciphertexts encrypted under distinct keys. Subsequent works by Chen and Dai et al. (CCS 2019) and Kim and Song et al. (CCS 2023) extended this concept by proposing multi-key BFV/CKKS variants, referred to as the CDKS scheme. These variants incorporate asymptotically optimal techniques to facilitate secure computation across multiple data providers. In this paper, we identify a critical security vulnerability in the CDKS scheme when applied to multiparty secure computation tasks, such as privacy-preserving federated learning (PPFL). In particular, we show that CDKS may inadvertently leak plaintext information from one party to others. To address this issue, we propose a new scheme, SMHE (Secure Multi-Key Homomorphic Encryption), which incorporates a novel masking mechanism into the multi-key BFV and CKKS frameworks to ensure that plaintexts remain confidential throughout the computation. We implement a PPFL application using SMHE and demonstrate that it provides significantly improved security with only a modest overhead in homomorphic evaluation. The code is publicly available at https://github.com/JiahuiWu2022/SMHE.git.
1 Introduction
@@
Homomorphic encryption (HE) is a cryptographic technique that enables computations to be performed directly on encrypted data, eliminating the need for decryption during the process. This property allows for the secure processing of sensitive data while preserving its confidentiality. For decades, constructing a HE scheme, capable of supporting arbitrary computations on ciphertexts, remained an open problem until Gentry’s groundbreaking work [1]. Since then, significant advancements have been achieved in the field, leading to various HE schemes such as BFV [2, 3], GSW [4], BGV [5], TFHE [6], and CKKS [7]. These schemes have expanded the practicality and efficiency of HE, making it a vital tool for secure computations in modern applications. One of the key features of HE is its ability to enable secure computation “on-the-fly,” meaning that data owners are not required to be actively involved during the computation process. Instead, the evaluation can be carried out entirely by a public server. This makes HE particularly appealing for scenarios such as cloud computing, where data is processed remotely, and privacy preservation is of paramount importance.
@@
In recent years, the demand for secure multiparty computation (MPC) protocols has surged, driven by applications such as federated learning (FL) [8]. MPC allows multiple parties to collaboratively evaluate a function or circuit on their private inputs without revealing any information beyond the final result. However, traditional single-key HE schemes are not well-suited for such multi-party settings. A major limitation arises when multiple data sources are involved, as standard HE schemes typically require all data to be encrypted under the same encryption key. This requirement grants the entity possessing the corresponding decryption key full access to the encrypted data, which raises issues regarding privacy and the reliance on centralized trust.
@@
In THE, multiple parties collaboratively generate a common public key, with the corresponding secret key shared among them through a secret-sharing mechanism. Although THE schemes generally achieve performance levels comparable to single-key HE and are typically more efficient than multi-key HE schemes, they suffer from a critical limitation: reliance on a static key access structure. In other words, all participants must be predetermined and fixed during the initial setup phase. In contrast, this work focuses on MKHE, which offers significant advantages in terms of flexibility and reduced interaction requirements. Specifically, MKHE schemes allow each participant to independently generate their own public-secret key pairs without needing knowledge of other parties’ keys. This property enables operations on ciphertexts encrypted under different keys and allows computations to be performed in a public cloud environment without establishing a common public key. Moreover, recent advances in MKHE have introduced full dynamism, allowing computations to be performed on multi-key ciphertexts without predefined circuits. Arbitrary circuits can be evaluated in real time, and new participants or ciphertexts can be incorporated into ongoing evaluations at any stage. This flexibility facilitates the construction of SMC protocols on top of MKHE, leveraging its dynamic and adaptive nature [17]. The ability to seamlessly integrate new participants and ciphertexts on-the-fly is a key advantage, making MKHE particularly well-suited for applications such as FL, where the set of participants may change throughout the computation process.
@@
1.1 Challenges and Contributions
Building on the above observation, we revisit the security assumptions behind CDKS-style MKHE and make the challenges and our remedies explicit. We keep technical details lightweight here and defer full derivations to later sections.
Challenge 1: CDKS violates the only-the-result goal in MPC. In the CDKS expansion + distributed-decryption workflow, the expanded component for party $i$ is tied to that party’s fresh share, and the merge uses the observable pair $(c_{i0},\nu_{i})$ (here $ct_{i}=(c_{i0},c_{i1})$ denotes a fresh BFV/CKKS ciphertext under key $i$, and $\nu_{i}$ is party $i$’s partial decryption share). Consequently, any party that sees $(c_{i0},\nu_{i})$ can recover client $i$’s plaintext via the one-line attack
@@ 1.1 Challenges and Contributions
contradicting the MPC ideal functionality that only the round result be revealed.
Remedy (structured masking). To eliminate this leakage channel, we attach to each input a structured masking triple $(ct,cz,\Gamma)$ and derive an auxiliary ciphertext $cx$ so that all intermediate states are statistically independent and become informative only at the final linear merge. Our design is tailored to CKKS/BFV and remains compatible with multi-key evaluation. Conceptually, masking is an encryption of fresh randomness that is added to the ciphertext; because the mask is itself a semantically secure ciphertext, intermediate views reveal nothing about the underlying message.
Provenance. The masking paradigm we adopt is not ad hoc: it adapts masking primitives originally developed for GSW-based multi-key HE (MKGSW) [16, 17] to the polynomial-ring setting of BFV/CKKS via a new ciphertext expansion and evaluation procedure. This preserves the only-the-result goal by removing the leakage path through $(c_{i0},\nu_{i})$ while sustaining efficient multi-key computation.
Challenge 2: Why MKGSW-style masking does not drop in for CKKS/BFV. While the MKGSW masking scheme ensures security and correctness in the GSW matrix setting, a direct transplant to CKKS/BFV fails for two reasons:
R1 (structural mismatch). GSW ciphertexts are matrices $C\in\mathbb{Z}_{Q}^{n\times m}$, whereas CKKS/BFV ciphertexts are polynomial pairs $ct=(c_{0},c_{1})\in R_{Q}^{2}$. This makes it nontrivial to realize a CKKS/BFV joint decryption relation of the form
@@
that mirrors the MKGSW equation $\big(sk(C+X)+sk^{\prime}C\big)\cdot\Theta^{-1}W\approx\mu$ (here $cx$ would be a mask derived from auxiliary info $\Gamma$, and $\theta$ and $\Theta$ are two fixed term).
R2 (quadratic blow-up). The GSW masking scheme requires large extended ciphertexts for correct demasking, with each single-bit encryption resulting in an $O(n^{2})$-size expanded ciphertext in the number of parties [^1] In the $n$-party setting, expansion realizes an $n\times n$ block layout: diagonal blocks carry per-party data while off-diagonal positions provide masking hooks to enable joint demasking. Both the row and column dimensions grow linearly with $n$, so the encoded length scales as $O(n^{2})$ times a base GSW block, even when many entries are structurally zero.. This leads to significant computational and communication overhead. Therefore, it is necessary to design a new masking scheme for CKKS/BFV that ensures the expanded ciphertext size scales linearly with $n$, thereby improving both computational and space efficiency.
Remedy (linear-size masked expansion). We propose a CKKS/BFV-specific masking scheme that achieves a linear-size expansion (i.e., the expanded ciphertext size scales linearly in $n$) and a correct joint decryption. In addition to a masking ciphertext $cx=(x_{0},x_{1})$ (carrying a random mask $r$), we generate a “demasking” ciphertext $cz=(z_{0},z_{1})$ and enforce
@@
Then the joint decryption of the expanded ciphertext
$\overline{ct}=(z_{0}+x_{0}+c_{0},\;x_{1}+c_{1},\;z_{1})\triangleq(\bar{c}_{0},\bar{c}_{1},\bar{c}_{2})$
under the concatenated key $\overline{sk}=(1,s,s^{\prime})$ yields
$\langle\overline{sk},\overline{ct}\rangle=\langle sk,ct\rangle+\underbrace{\big(\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\big)}_{\approx 0}\approx\mu,$
and the size of $\overline{ct}$ is linear in the number of participants.
Instantiation (CKKS; BFV is analogous). Let $ct\leftarrow\texttt{Encrypt}(pk,\mu)=w\cdot pk+(\mu+e_{0},e_{1})\bmod Q$, where $sk=(1,s)$, pk=(b=as+e,a)pk=(b=-as+e,\,a), and $e,e_{0},e_{1}$ are small. Choose $cz\leftarrow\texttt{Encrypt}(pk,0)=r\cdot pk+(e_{r_{0}},e_{r_{1}})$ so that
$\langle sk^{\prime},cz\rangle=r(b-b^{\prime})+e_{z}\bmod Q,\quad b^{\prime}=-as^{\prime}+e^{\prime}\bmod Q.$
Encrypt the mask $r$ as auxiliary information $\Gamma\leftarrow\texttt{Encrypt}(pk,r)=w\cdot pk+(r+e_{x_{0}},e_{x_{1}})$, and set $cx\;=\;\Gamma\,(b^{\prime}-b),$ so that, within the decryption noise margin, $\langle sk,cx\rangle$ cancels $\langle sk^{\prime},cz\rangle$ and thereby satisfies Eq. (1).
$n$-party generalization (linear growth). For $n$ parties with secret keys $sk_{i}=(1,s_{i})$, we create per-party masked pairs $(cx_{i},cz_{i})$ and aggregate them into a single expanded ciphertext whose coordinate-length increases by exactly one for each additional party:
$\overline{ct}=(\bar{c}_{0},\bar{c}_{1},\ldots,\bar{c}_{n}),\qquad\overline{sk}=(1,s_{1},\ldots,s_{n}).$
The per-party cancellation constraints (generalizing Eq. (1)) ensure that mask contributions vanish in the final inner product, giving the unified joint decryption
$\langle\overline{sk},\overline{ct}\rangle=\big\langle(1,s_{1},\ldots,s_{n}),\,(\bar{c}_{0},\ldots,\bar{c}_{n})\big\rangle\approx\mu.$ (2)
Hence the expanded ciphertext has (n+1)(n{+}1) coordinates and its size scales linearly in $n$.
Challenge 3: Noise growth breaks naive demasking. The naive choice $cx=\Gamma\,(b^{\prime}-b)$ introduces a large error term due to multiplication by the essentially random polynomial $(b^{\prime}-b)\in R_{Q}$:
@@
$=r(b^{\prime}-b)+\underbrace{(b^{\prime}-b)(we+e_{x_{0}}+e_{x_{1}}s)}_{E}\;\bmod\ Q,$
where $E\in R_{Q}$ can be non-negligible. Consequently, the desired cancellation $\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\approx 0$ fails, invalidating the decryption equation.
Remedy (Gadget decomposition to tame noise). We control the error growth by applying gadget decomposition to the factors that amplify noise, reducing $E$ so that $\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\approx 0$ holds with the necessary tightness. Intuitively, gadget techniques linearize large multipliers into well-behaved digit expansions that keep the induced error within the decryption margin. We provide the decomposition details in Section 3.3 and the full masking construction in Section 5.1.
2 Related Work